A crane used 112.95 watts of power to raise a 17.74 N object in 4.05 seconds. Through what vertical distance was the object displaced?





1
Expert's answer
2022-02-11T08:47:08-0500

Given:

"P=112.95\;\rm W"

"F=17.74\;\rm N"

"t=4.05\:\rm s"


The work done by the crane

"W=Fh=Pt"

Hence, the vertical distance the object was displaced

"h=\frac{Pt}{F}=\rm\frac{112.95\: W*4.05\: s}{17.74\: N}=25.8\:\rm m"